An analytical chemist is titrating
91.1mL
of a
0.9300M
solution of propionic acid
HC2H5CO2
with a
0.2700M
solution of
KOH
. The
pKa
of propionic acid is
4.89
. Calculate the pH of the acid solution after the chemist has added
346.mL
of the
KOH
solution to it.
1
Expert's answer
2017-03-28T02:11:07-0400
The chemical equation is
CH3-CH2-COOH + KOH → CH3-CH2-COOK + H2O
Acid____________base

Using given data we can determine tne number of moles for each reactants:
n(acid) = C(acid)∙V(acid) = 0.9300∙0.09110 = 0.08472 (mol)
n(base) = C(base)∙V(base) = 0.2700∙0.3460 = 0.09342 (mol) = n(CH3-CH2-COOK)
In equilibrium state we have buffer system as CH3-CH2-COOH + CH3-CH2-COOK. Thus, we can calculate the pH value of this buffer solution, using formulae:
pH = pKa + lg(n (CH3CH2COOK))/(n(CH3CH2COOH)) = 4.89 + lg 0.09342/0.08472 = 4.93
If pH = 4.93<7, so this solution is acidic.
Answer: pH = 4.93.